# Relative Camera Pose Regression

### v1

@@ 6 Conclusion and Future Work
In this paper, we introduced Alligat0R, a novel pre-training approach for binocular vision that reformulates cross-view completion as a co-visibility segmentation task. Unlike CroCo that relies on masked reconstruction, our approach explicitly models co-visibility relationships between image pairs, enabling more effective learning from challenging examples with limited overlap.
We demonstrated the effectiveness of our approach on the relative pose regression task in the autonomous driving scenario, where Alligat0R significantly outperforms CroCo, particularly on challenging scenes with limited overlap between views. This improvement is especially notable using a diverse training set that include pairs with varying degrees of overlap, highlighting the ability of our approach to learn and benefit from geometrically challenging examples.
To enable our approach, we created Cub3, a large-scale dataset of 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset. This dataset provides a valuable resource for future research in binocular vision tasks and will be made publicly available.
Our work demonstrates that explicitly modeling co-visibility through segmentation provides a more effective and interpretable approach to binocular vision tasks compared to reconstruction-based methods, paving the way for more robust visual perception systems.
Supplementary Material
7 Supplementary Material
@@
In this supplementary material, we provide additional details about our Alligat0R model, including training curves and visualizations that didn’t fit in the main paper but may be of interest to readers.
7.1 Pre-training Learning Curves
@@
Figure 7 shows the learning curves for the pre-training phase of both CroCo and Alligat0R on the Cub3-50 and Cub3-all datasets. As can be observed, Alligat0R’s loss converges smoothly on both datasets, indicating stable training. There are fewer number of steps for CroCo as we have a batch size of 64 per GPU, compared to Alligat0R where we have a batch size of 32.
@@
Figure 7: Learning curves during pre-training for CroCo (reconstruction loss) and Alligat0R (segmentation loss) on Cub3-50 and Cub3-all datasets. Alligat0R shows stable convergence on both datasets, while CroCo exhibits instability on the more challenging Cub3-all dataset which contains image pairs with lower overlap.
7.2 Fine-tuning Learning Curves
@@
Figure 8 presents the fine-tuning curves for the pose regression task. The plots show the pose loss during training. Alligat0R fine-tuned on Cub3-all shows faster convergence and reaches higher accuracy than the same configuration for CroCo, highlighting the transferability of features learned through co-visibility segmentation on difficult pairs.
@@
Figure 8: Learning curves during fine-tuning for pose regression. Alligat0R pre-trained on Cub3-all converges faster and achieves higher success rates than other methods, demonstrating the effectiveness of our co-visibility segmentation pre-training approach.
7.3 Detailed Performance
@@
Figure 9 provides a more detailed breakdown of performance across different geometric criteria than what was presented in the main paper. This visualization further emphasizes Alligat0R’s strong performance across all difficulty ranges, especially on very difficult scenarios where traditional methods struggle.
@@
Figure 9: Detailed breakdown of performance across different geometric criteria. Success rate either for R@5° or t@2m (bottom-left and top-right of each triangle, respectively), when projecting results onto individual geometric criteria of RUBIK. For each method, we show three plots corresponding to the projection over overlap (left), scale ratio (middle), and viewpoint angle (right)

### v2

@@ 7 Conclusion and Future Work
We introduced Alligat0R, a novel pre-training approach that replaces cross-view completion with covisibility segmentation. Unlike CroCo’s masked reconstruction, our approach explicitly models covisibility relationships, enabling more effective learning from challenging geometric configurations.
We demonstrated significant improvements over CroCo on metric relative pose regression and zero-shot correspondence estimation, particularly on challenging scenes with limited overlap. Our approach’s robustness to geometrically challenging examples and superior generalization capabilities highlight the effectiveness of covisibility segmentation for binocular vision tasks.
We created Cub3, a large-scale dataset of 5M image pairs with dense covisibility annotations from nuScenes and ScanNet. Our work demonstrates that explicitly modeling covisibility provides an effective alternative to binocular vision tasks compared to reconstruction-based methods.
Acknowledgments
This project has received funding from the Bosch Research Foundation (Bosch Forschungsstiftung), the European Union (ERC Advanced Grant explorer Funding ID #101097259), and was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014905R1 made by GENCI.
Appendix A Supplementary Material
@@
This supplementary material provides additional details about our Alligat0R model, including training curves, implementation details, and visualizations that complement the main paper.
A.1 Pre-training Learning Curves
@@
Figure 7 shows the learning curves for the pre-training phase of both CroCo and Alligat0R on the Cub3-50 and Cub3-all datasets. Alligat0R’s loss converges smoothly on both datasets, indicating stable training.
@@
Figure 7: Learning curves during pre-training for CroCo (reconstruction loss) and Alligat0R (segmentation loss) on Cub3-50 and Cub3-all datasets for nuScenes (left) and ScanNet (right). Alligat0R shows stable convergence on both datasets.
A.2 Fine-tuning Learning Curves
@@
Figure 8 presents the fine-tuning curves for the pose regression task. The plots show the full loss during training (Eq. 9 in the main paper for Alligat0R, Eq. 8 for CroCo). Alligat0R fine-tuned on Cub3-all shows faster convergence and reaches higher accuracy than the same configuration for CroCo, highlighting the transferability of features learned through covisibility segmentation on difficult pairs. The loss increase at 5 epochs corresponds to unfreezing the backbone. For Alligat0R, at 5 epochs we also add the segmentation loss to maintain interpretability.
@@
Figure 8: Learning curves during fine-tuning for pose regression for nuScenes (left) and ScanNet (right). Alligat0R pre-trained on Cub3-all converges faster and achieves higher success rates than other methods, demonstrating the effectiveness of our covisibility segmentation pre-training approach.
A.3 Detailed Performance Analysis
@@
Figure 9 provides a comprehensive breakdown of performance across different geometric criteria on RUBIK. This visualization emphasizes Alligat0R’s strong performance across all difficulty ranges, particularly in challenging scenarios where traditional methods struggle.
@@
Figure 9: Detailed breakdown of performance across different geometric criteria. Success rate either for R@5° or t@2m (bottom-left and top-right of each triangle, respectively), when projecting results onto individual geometric criteria of RUBIK. For each method, we show three plots corresponding to the projection over overlap (left), scale ratio (middle), and viewpoint angle (right).
A.4 Implementation Details
We use a ViT-based encoder and transformer decoder backbone similar to CroCo, with 24 layers for the encoder and 12 for the decoder. For pre-training, we use the AdamW optimizer with a learning rate of 1.5e-4, weight decay of 0.05, and a batch size of 32 per GPU. We employ a cosine learning rate schedule with 2 epochs of warmup and train for 25 epochs on our Cub3 datasets.
For fine-tuning on relative pose regression, we follow the two-phase approach described in Section 3.3 of the main paper:
  • Phase 1: Freeze the backbone and train only the pose regression head for 5 epochs with a learning rate of 1e-4
  • Phase 2: Unfreeze the entire network and jointly train with both the pose regression loss and covisibility segmentation loss for an additional 10 epochs with a learning rate of 5e-5
The comprehensive architecture of Alligat0R is illustrated in Figure 2 of the main paper, showing both the pre-training and fine-tuning phases. During pre-training, the model learns to segment pixels in each view as covisible, occluded, or outside the field of view with respect to the other view. During fine-tuning, we introduce a pose regression head while maintaining the segmentation capability to leverage the geometric understanding acquired during pre-training.
A.5 Data Splits and Sampling Protocol
To ensure proper evaluation and avoid data leakage, we carefully separate training and test data:
Data Splits:
  • RUBIK benchmark: Uses test scenes from nuScenes, while our Cub3 dataset uses only the training split of nuScenes, with completely disjoint scenes.
  • ScanNet1500 benchmark: Derived from the standard ScanNet test split, whereas Cub3 uses scenes from the ScanNet training split exclusively.
Sampling Protocol: We pre-filter all possible pairs from all training scenes with at least 5% overlap for the "all" version of Cub3, and at least 50% for the "50" version. We then sample from all those pre-filtered pairs to extract the desired number of samples and covisibility masks. The RUBIK benchmark creation follows the protocol described in [loiseau2025rubik] and is extracted from test scenes, while ScanNet1500 was created from test scenes of ScanNet as first introduced in [sarlin2020superglue].
A.6 Additional Ablation Studies
A.6.1 Impact of the Number of Classes
We implemented two variants of Alligat0R to investigate the importance of our pre-training with the three classes (covisible, occluded, outside FOV). We tried pre-training with only two classes, either covisible or not (in this case occluded and outside FOV are merged), or inside FOV or not (in this case, covisible and occluded are merged) on Cub3-all for nuScenes. The results on the RUBIK benchmark are presented in Table˜5.
Table 5: Results on RUBIK for metric relative pose regression when pre-training with only two classes. For all experiments, pre-training and fine-tuning is performed using Cub3-all.
RUBIK
Classes $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m
Covisible or not 22.5 55.7 80.0
Inside FOV or not 24.6 59.6 81.9
All 3 classes 24.6 60.3 81.9
While the performance improvement with three classes is modest, we believe that the model’s knowledge about occluded regions could be beneficial for other tasks. As noted in the main paper, the annotated maps may contain noise between covisible and occluded zones due to reliance on monocular depth predictions.
A.6.2 Non-metric Relative Pose Regression
We investigated whether our pre-training is useful for non-metric relative pose regression (regressing only the angle for translation) and compared our results with CroCo pre-training on the ScanNet1500 benchmark. For this experiment, we changed our pose regression head using the one from Reloc3r, along with its loss function. The results are shown in Table˜6.
Table 6: Results on ScanNet1500 for non-metric relative pose regression. For Alligat0R, pre-training and fine-tuning is performed using Cub3-all, whereas for CroCo, pre-training is performed using Cub3-50 and fine-tuning using Cub3-all.
ScanNet1500
Pre-Training AUC@5 AUC@10 AUC@20
CroCo 13.2 34.1 57.1
Alligat0R 20.5 43.9 66.2
Alligat0R significantly outperforms CroCo on this task, demonstrating the versatility of our pre-training approach.
A.7 Zero-shot Correspondence Estimation on ETH3D
To demonstrate the generalization capabilities of our method beyond the training domains, we evaluate Alligat0R on zero-shot correspondence estimation using the ETH3D dataset [schops2017multi]. We use the correlations from decoder features to estimate correspondences and measure performance using Average EndPoint Error (AEPE) as described in [an2025cross].
Table 7: Zero-shot correspondence estimation on ETH3D dataset using AEPE. Alligat0R consistently outperforms CroCo variants, demonstrating better generalization to out-of-domain dense matching tasks.
Method Training Data AEPE ($\downarrow$)
Alligat0R nuScenes-all 43.82
CroCo nuScenes-all 77.61
Alligat0R nuScenes-50 44.12
CroCo nuScenes-50 92.98
Alligat0R ScanNet-all 36.07
CroCo ScanNet-all 56.70
Alligat0R ScanNet-50 38.45
CroCo ScanNet-50 86.60
CroCo v2 5 datasets 51.55
The results in Table˜7 show that Alligat0R consistently outperforms CroCo when trained on the same datasets, and even surpasses CroCo v2 which was pre-trained on 5 datasets. This demonstrates that our covisibility segmentation pre-training learns more generalizable features for dense correspondence tasks.
A.8 Additional Qualitative Results
We provide additional visualizations from our nuScenes and ScanNet datasets in Figures˜10 and 11, along with more predictions from Alligat0R and CroCo in Figure˜12.
Figure 10: Covisibility annotation examples from Cub3 for nuScenes. For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions. Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes. Let us highlight that some annotations, particularly the distinction between covisible and occluded pixels, may contain noise, especially for nuScenes, and we demonstrate in the experiments that Alligat0R is highly robust to this noise.
Figure 11: Covisibility annotation examples from Cub3 for ScanNet. For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions. Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes.
 Reference    Target       Masked   CroCo recons.   Alligat0R segmentations
Figure 12: Qualitative comparison. CroCo’s reconstructions are blurred in masked non-covisible regions, while Alligat0R often correctly identifies covisible, occluded, outside FOV regions across varying degrees of overlap and viewpoint changes.